OpenStudy (anonymous):

Σ (1 + sin n)/10^n from 0 to infinity, determine the convergence...Can someone work out the answer? I'm completely lost on how to figure these out..AWARD AT END!

OpenStudy (mathmate):

hmm

OpenStudy (anonymous):

@mathmate Got any ideas?

OpenStudy (mathmate):

You realize that by using the sandwich theorem, the sum is between 0/10^n and 2/10^n. This way, you can get rid of the ugly sine term.

OpenStudy (mathmate):

After that, you only have a GP to worry about, and the sum (of the limits) are not difficult to calculate.

OpenStudy (anonymous):

Well wouldn't 0/10^n just be 0? And no I wouldn't know that...That's why I am asking for help

OpenStudy (mathmate):

The sum of 0/10^n would evidently be zero, and that's the lower bound.

OpenStudy (anonymous):

Okay the convergence is 20/9 then. It converges

OpenStudy (anonymous):

Thanks for the help

OpenStudy (mathmate):

It's the 2/10^n that you need to calculate as the sum of a geometric progression, with common ration 10.

OpenStudy (mathmate):

*ratio

OpenStudy (anonymous):

Do you think you could help me solve the integral of convergence of sqrt(1+x)...I'm studying these for my class this week and I can't figure them out.

OpenStudy (mathmate):

I can try, if you post the question. Not sure if I can do it.

OpenStudy (anonymous):

Well that is the question. Find the power series and interval of convergence of sqrt(1+x). I already used the Taylor Series to get the "power series" but I can't get the interval of convergence. I keep on getting lost in the steps. I need someone to show me them.

OpenStudy (anonymous):

Use integral test: [\int\limits_{1}^{\infty} \frac{ 1 }{10 ^{x}}+\frac{ \sin x }{10 ^{x} }dx\] You will find that it tends to .0827487 (Wolfram Alpha, or you can integrate the terms by hand). According to the integral test if the integral is convergent then \[\sum_{1}^{\infty} \frac{ 1+\sin n }{ 10^{n} }\] is also convergent. Therefore \[\sum_{0}^{\infty}\frac{ 1+\sin x }{10^{x} }\] is also convergent because when x=0, the summand isfinite.

OpenStudy (anonymous):

Thank you @DominicNg :) That really does help. Wolfram Alpha is a great math tool! Wish they didn't charge for it. -_-

OpenStudy (mathmate):

@dominicNg It is better than you think, because the summand is not infinite when x=0. 10^0 turns out to be 1. :)

OpenStudy (anonymous):

It is convergent for any n.

OpenStudy (mathmate):

@zerosniper123 So did you expand sqrt(1+x), and what did you get.

OpenStudy (mathmate):

The interval of convergence is the range of x for which the series will converge.

OpenStudy (anonymous):

I don't know why I can't paste it...I used a math program to write it out...Im just gonna a pic of it.

OpenStudy (anonymous):

@mathmate when I expanded it and found the power series using Taylors method I got that.

OpenStudy (mathmate):

You need to show convergence for the maximum range of x. Since it is an alternating series, convergence is easier to prove.

OpenStudy (anonymous):

Okay, well could you do the steps?? Right now I am brain fried...I don't really get this stuff...I understand taylor series and power series and Maclaurin...But trying to take the integral.

OpenStudy (mathmate):

The series you gave me is correctly expanded. This is the very first step.

OpenStudy (anonymous):

Whats the next step?

OpenStudy (mathmate):

Are you asked to find the interval of convergence using the integral test?

OpenStudy (anonymous):

Its not asking me to use the integral test, just find the interval of convergence. @mathmate

OpenStudy (anonymous):

Radius of convergence is 1. Interval of convergence is (-1,1). See enclosed.

OpenStudy (mathmate):

@DominicNg thank you for the detailed solution. I hope you wouldn't mind if @zerosniper123 would ask for explanations on the solution.

OpenStudy (anonymous):

What would you like to know